A 11.0 cm wire is being pulled along a U-shaped wire at a speed of 8.20 m/s. The system is in a perpendicular magnetic field. The total resistance of the system is 0.100 Ω. If the power dissipated in the system is 5.10 W, what is the strength of the magnetic field?


A. 0.714 T
B. 5.05 T
C. 0.00565 T
D. 0.792 T
Answer #1

Power P = VI = V^2/R = (BvL)^2/R

=) B = 1/vL × √PR = √(5.10×0.1) / 8.2×0.11 = 0.792T

=) B = 0.792 T ; option D